# Does excellence correspond to universal inequality level? Evidences from scholarly citations & Olympic medal data

### v2

@@ III.1 Inequality in citations
However, a more standardized measure for the segregation of the Nobel laureates would be to construct a Receiver Operation Characteristic (ROC) curve Tanner and Swets (1954) for each of these indices (or combinations thereof), such that an objective quantification of their relative success in distinguishing the Nobel laureates could be compared. In general, an ROC curve is a plot of True Positive Rate (TPR) with the False Positive Rate (FPR) in an attempted segregation as the parameter in question is gradually varied (see Fawcett (2006) for a review). In the present context, suppose we want to quantify the usefulness of any quantity $J$ in doing the classification. We then measure the average of that quantity for all Nobel laureates. Then we consider the standard deviation of $J$ among the Nobel laureates. Then, we calculate the ratio of the fraction of Nobel laureates within the range $\langle J\rangle-\Delta J$ to $\langle J\rangle+\Delta J$, and we calculate the ratio of other scientists, within the exact same range, where $\Delta J=n\sigma$, with $\sigma$ being the standard deviation of $J$ and $n$ is a number that we gradually increase. The fraction of Nobel laureates is then the TPR and the fraction of others is the FPR and we plot the points calculated for different values of $n$ (see Fig. 6), giving us the ROC curve. The ROC curve, by definition, is bounded between (0,0) and (1,1). Clearly, a 45-degree line is when the classifier works no better than random choices. The area under the curve (AUC) between the ROC and 45-degree line is then a measure of the performance of the classifier. This is then a standardized measure for any quantity, particularly the inequality indices, even though the quantities individually may have different ranges of their own. In this way we can have a quantification of the comparison of performances of different scientometric metrices in segregating the Nobel laureates.
In what follows, we mostly focus on the two indices $g$ and $k$ and leave out $h$ and $Q$ due to their unbounded nature, as these could significantly vary post-award and thereby places the different laureates in different footings depending on the time elapsed since their award until now. As shown in Fig. 6, we compare three different quantities: $g$, $k$, and their ratio $g/k$. Each of these measures serves as a relatively good indicator of citation inequality. However, the selection of a specific metric depends on the acceptable trade-off between false and true positives. Notably, $g/k$ appears to perform slightly better than $g$ and $k$ individually, suggesting that the combined measure may capture inequality patterns more effectively.
III.2 Inequality in Olympic medals
Finally, we turn to the question of inequality of Olympic medals received by various countries in the summer and winter versions of the games over the period of 1896-2024 and 1922-2022 respectively. As mentioned before, in this case also, a high level of competition is present. In the similar way as described before, one can construct the Lorenz curve and then calculate the inequality indices $g$ and $k$ for the medals received by different countries in a given year.
In Fig. 7, the time variation of $g$ and $k$ are shown for the summer Olympic medals. Note that the number of participating countries have changed drastically over the years (see Table 2 in the Appendix). In more recent times, the inequality indices $g$ and $k$ have remained close to each other and around $0.85$. A similar trend can also be observed for the winter Olympic games (see Table 3 in the Appendix) as well.
@@ IV Discussions and conclusions
Reduction of inequality is one of Sustainable Development Goals by the United Nations since 2015 United Nations SDGs (2023). However, a little more than Pareto’s 80-20 (or Self-organised Critical) level of social inequality seems inherently tied to social competence and efficiency. An interesting question would be how far such high level inequalities can be sustained in different spheres of social dynamics? Such questions are addressable, particularly in the contexts where accumulated asset inequalities are not tied to any dire consequences, or even sometimes a desirable phenomenon. If a steady state then can emerge out of the collective behavior of the competing individuals, it is then useful to inquire what are the characteristics of such agents who accumulates a higher amount of asset.
@@
In this work, we tried to proceed along that direction, particularly in the context of scholarly citations among the papers of individual researchers and the Olympic medals (summer and winter) received by different countries in a year. The two situations are slightly different from one another. In the first case i.e., the inequality of scholarly citations of different papers of many researchers, we note that researchers who are widely recognized as successful (Nobel laureates) have a higher inequality in the citations of their individual papers. More particularly, the Nobel laureates are disproportionately concentrated near the higher values of the inequality indices (Gini $g$ and Kolkata $k$ indices) when labeled with the inequality indices associated with them. We have made a quantitative analysis regarding the effectiveness of using the $g,k$ and $g/k$ indices in distinguishing Nobel laureates from the other scientists (see Fig. 6). It then suggests that inequality in this context could be a useful indicator of excellence. Here, we focus on the bounded indices ($g\in(0,1)$, $k\in(0.5,1)$, and $g/k\in(0,2)$) rather than unbounded ones such as the Hirsch index or Q-factor ($h,Q\in(0,\infty)$), as the latter, despite potentially offering higher separation power based on ROC analysis, may be less reliable. Post-award, these unbounded indices can be disproportionately affected by the scientist’s recognition, whereas bounded measures may remain more stable. Further quantitative analysis is required to validate this observation by examining the time evolution of these indicators and comparing pre- and post-award behavior.
@@
We then also look at the case of inequality in winning the Olympic medals (both in the summer and winter Olympics) among the participating countries. Note that Olympic games are already the most competitive athletic event. Therefore, arguing along the same line as above, one would expect the medal distribution to be highly unequal, which is exactly what we find, especially when the number of participating countries are large. The data for the Olympic medal inequalities should then be compared with the data points for the Nobel laureates in the citation case, and the data for the other scientists would have been similar to inequality in the medal tally of some other games, which is not as competitive as the Olympics. It is hard to conceive of such an example in the similar scale of participation and we did not investigate along that line.
@@
Finally, a couple of more points need to be noted for the citation analysis. Firstly, the number of Nobel laureates is restricted by other conditions than just excellence. There are only a fixed maximum number of laureates in a field in year, this is not awarded posthumously and so on. This means, there could be many scientists in the range of citation inequality as that of the Nobel laureates and as excellent. Therefore, the False Positive we mentioned earlier, is not necessarily ’false’ in terms of distinguishing excellent researchers. A more detailed analysis for the scientists who fall in this range could be a fruitful future direction of research. Finally, an intuitive understanding as to why such citation inequalities are seen for Nobel laureates could be argued in the following way: If many papers of a researcher receive similar level of (high) citations, a more probable explanation for that could simply be the higher rate of publications in their particular field rather than all those papers being outstanding. This situation would result in a high $h$-index but low $g,k$ and $Q$. We don’t see that for Nobel laureates. In this case, the Nobel winning works would have received much more attention (even prior to the award) than other works of the laureate. There could of course be individual exceptions. The question of not observing extreme inequality ($g,k$ near unity) for Nobel laureates is more subtle. It could be expected that except for the Nobel winning work, an outstanding researcher such as a Nobel laureate would have made some other important contributions as well that would receive good citations and thereby preventing the inequality to reach an extreme level.
@@
In conclusion, we find emergent inequality is a good indicator of high competitiveness and excellence. We demonstrate this through extensive data analysis for scholarly citations and Olympic medal tallies over the years.

### v3

@@ IV Discussions and Conclusions
Reduction in inequality is one of the Sustainable Development Goals that has been established by the United Nations since 2015 United Nations SDGs (2023). However, a little more than Pareto’s 80–20 (or self-organized critical) level of social inequality seems inherently tied to social competence and efficiency. An interesting question would be how far such high-level inequalities can be sustained in different spheres of social dynamics? Such questions are addressable, particularly in the contexts where accumulated asset inequalities are not tied to any dire consequences, or even sometimes a desirable phenomenon. If a steady state can then emerge out of the collective behavior of the competing individuals, it is then useful to inquire what the characteristics of such agents who accumulate a higher amount of assets are.
@@
In this work, we tried to proceed along that direction, particularly in the context of scholarly citations among the papers of individual researchers and the Olympic medals (summer and winter) received by different countries in a year. The two situations are slightly different from one another. In the first case, i.e., the inequality of scholarly citations of different papers of many researchers, we note that researchers who are widely recognized as successful (Nobel laureates) have a higher inequality in the citations of their individual papers. More particularly, the Nobel laureates are disproportionately concentrated near the higher values of the inequality indices (Gini $g$ and Kolkata $k$ indices) when labeled with the inequality indices associated with them. We have made a quantitative analysis regarding the effectiveness of using the $g,k$ and $g/k$ indices in distinguishing Nobel laureates from the other scientists (see Figure 7). We have also studied the effectiveness, using the ROC and variants of it, for several other measures, such as the entropy $S$, TP parameter $b$, $Q$ factor and Hirsch index $h$. It then suggests that inequality in this context could be a useful indicator of excellence. This is in line with the observation that scientific impact likely requires more than one index for quantification 3d_index ; born .
As mentioned before, the parameters we study can be divided broadly into three groups depending on their domain of values, bounded ($g$, $k$), semi-bounded ($S$, $b$) and unbounded ($Q,h$). Post-award, the unbounded indices can be disproportionately affected by the scientist’s recognition, whereas bounded and semi-bounded measures may remain more stable. Further quantitative analysis is required to validate this observation by examining the time evolution of these indicators and comparing pre- and post-award behavior. We also fitted the citation count distributions of individual researchers using the mean-normalized Tsallis–Pareto (Lomax II) distribution. A clear distinction was observed in the fitted shape parameter $b$, with Nobel laureates consistently exhibiting values closer to the lower limit $b\to 1$. This suggests that they lie near the boundary of typical citation patterns, in line with their position at the frontier of scientific excellence. Additionally, the Shannon differential entropy computed from normalized citation distributions was found to be systematically lower for Nobel laureates, indicating a more ordered structure in their citation profiles.
It would also be valuable to explore the entropic properties of the Kolkata index, particularly through the maximum of gintropy, and examine its connection to the differences between Nobel laureates and other scientists.
@@
We then also look at the case of inequality in winning the Olympic medals (both in the Summer and Winter Olympics) among the participating countries. Note that the Olympic Games is already the most competitive athletic event. Therefore, arguing along the same lines as above, one would expect the medal distribution to be highly unequal, which is exactly what we find, especially when the number of participating countries is large. The data on the Olympic medal inequalities should then be compared with the data points for the Nobel laureates in the citation case. In contrast, the data for the other scientists may resemble the medal tally of less-competitive sporting events, where inequality is not as pronounced as in the Olympics. It is difficult to conceive of such an example in a similar scale of participation, and we did not investigate along this line.
@@
Finally, a couple of more points need to be noted for the citation analysis. Firstly, the number of Nobel laureates is restricted by other conditions than just excellence. There are only a fixed maximum number of laureates in a field in a year, which is not awarded posthumously, among other factors. This means that there could be many scientists in the range of citation inequality as that of the Nobel laureates who are just as excellent. Therefore, the false positive we mentioned earlier is not necessarily “false” in terms of distinguishing excellent researchers. A more detailed analysis of the scientists who fall in this range could be a fruitful future direction of research. Finally, an intuitive understanding as to why such citation inequalities are seen for Nobel laureates could be argued in the following way: If many papers of a researcher receive a similar level of (high) citations, a more probable explanation for that could simply be the higher rate of publications in their particular field rather than all those papers being outstanding. This situation would result in a high $h$-index but low $g,k$ and $Q$. We do not, however, see this for Nobel laureates. In this case, the works awarded Nobel prizes would have received much more attention (even prior to the award) than other works of the laureates. There could of course be individual exceptions. The question of not observing extreme inequality ($g,k$ near unity) for Nobel laureates is more subtle. It could be expected that except for the Nobel-winning work, an outstanding researcher such as a Nobel laureate would have made some other important contributions as well that would receive good citations, thereby preventing the inequality to reach an extreme level.
@@
In conclusion, we find that emergent inequality is a good indicator of high competitiveness and excellence. We demonstrate this through extensive data analysis for scholarly citations and Olympic medal tallies over the years.